OpenStudy (anonymous):

Given f(x) = x^ 2+3 and g(x) = x+5/x find (g o f)(4)

OpenStudy (anonymous):

I got 14/19

OpenStudy (anonymous):

@undeadknight26

undeadknight26 (undeadknight26):

I hate thees

OpenStudy (anonymous):

I hate them as much as you do, and it's barely the second day of school, sheesh.

undeadknight26 (undeadknight26):

Sorry m8 @dumbcow @Koikkara @Lady.Liv1776 @LunyMoony @whpalmer4 @rishavraj @rvc

OpenStudy (anonymous):

. . . xD

OpenStudy (lunymoony):

What does the 'o' mean in " (g o f)" first though?

OpenStudy (anonymous):

http://prntscr.com/88qogy

OpenStudy (dumbcow):

i dont like the notation for these (f o g) is same as f(g) so (g o f)(4) = g(f(4))

rvc (rvc):

g of f

OpenStudy (lunymoony):

Ok, thx

OpenStudy (dumbcow):

plug in 4 into "f", then plug that num into "g"

OpenStudy (anonymous):

f(4) = x^2+3 and g(4)?

OpenStudy (dumbcow):

no , finish f(4) first .... what do you get?

OpenStudy (anonymous):

Finish f(4)? How do I do that?

OpenStudy (dumbcow):

\[f(x) = x^2 +3\] \[f(4) = 4^2 + 3\] --you substitute the "x" with "4"

OpenStudy (anonymous):

Oh I figured. f(4)=4^2+3 g(4)(4)?

OpenStudy (anonymous):

Nevermind I don't know what I'm doing.

OpenStudy (anonymous):

Can someone just tell me if I'm right or not? xD

OpenStudy (dumbcow):

oh sorry, so whats 4^2 +3 ?

OpenStudy (anonymous):

Oh I thought that was the final product. I don't know f(x)

OpenStudy (dumbcow):

you are over thinking it..... 4^2 +3 = what

OpenStudy (anonymous):

19?

OpenStudy (dumbcow):

very good.... now use this number for function g ---> g(19) = (x+5)/x = (19+5)/19 = ?

OpenStudy (anonymous):

1.26?

OpenStudy (dumbcow):

yeah or just 24/19 \[g(f(4)) = g(19) = \frac{24}{19}\]

OpenStudy (anonymous):

Oh, thank you so much!

OpenStudy (dumbcow):

yw